# Are You Doubtful? Oh, It Might Be Difficult Then! Exploring the Use of Model Uncertainty for Question Difficulty Estimation

### v1

@@ 4.2 Model Uncertainty
The current methodological approach is founded on the premise that model uncertainty correlates with student performance and thus, by extension, offers a useful signal when estimating the difficulty of a question item. To explore this hypothesis, we have conducted experiments using two metrics that are shown to correlate well with model correctness (as discussed in Appendix  A): 1st Token Probability and Choice Order Sensitivity. These uncertainty scores are obtained for each LLM separately and concatenated into a single vector, to which textual features are (optionally) also added. This vector is then fed to the regressor.
@@
The first technique to measure model uncertainty is by inspecting the softmax probability of the 1st token to be generated as the answer id to the given MCQ question, (e.g., probability of generating token "B"), in comparison to the probabilities of the alternatives (e.g., probability of generating token "A" or "C"). As the 1st token probabilities can be influenced by the order in which the choices are provided in the problem set Wei et al. (2024); Wang et al. (2023, 2024); Zheng et al. (2024), we create ten random different orderings for each question and let the model answer each MCQ ten times [^6] For questions with only 3 choices, we instead consider all six different choice orderings.. This way, we calculate the average probability per MCQ choice. We then consider the average probability for the correct answer as the uncertainty metric of the LLM.
Furthermore, as different tokens might be generated to represent the same answer (e.g., "A", " A", "a ", see details on prompting and answer elicitation in Section 4.3 below) and different models might attribute higher likelihood for specific tokens, the token representing each choice with the highest probability is selected. For example if for a given model the probability of generating token "C" is higher than the probability of token "c", the former is considered for that model. Lastly, the three extracted mean probabilities of all orderings are normalised in the range of 01010-10 - 1 such that they can be more easily compared to the difficulty scores, which – being calculated as proportions of the student populations – are in the same range.
@@
Pezeshkpour and Hruschka (2023) observed that choice order sensitivity correlates with error rate. In other words, when LLMs consistently select a choice regardless of its position, that choice is more likely to be correct. Based on this observation, we leverage this correlation to measure uncertainty. Specifically, for all evaluated choice orderings, we measure the probability of the correct choice being selected. Thus, this probability is not based on token probabilities but rather on the eventual choice.
@@ 4.3 Choice of Models and Prompting
In this work, we focus on decoder-only models, as they are considered to have incorporated greater amounts of factual knowledge as a byproduct of their language modelling objective Zhao et al. (2023), compared to Encoder-Only or Encoder-Decoder models. Moreover, as the internal logit probabilities of the 1st token to be generated are needed to measure the uncertainty of each model, we focus on nine open-sourced models of different parameter sizes and families. Additionally, we constrict our choice to instruction-tuned models and use 4-bit quantisation for increased efficiency. [^7] Details of the models used are in Table 6 in Appendix C. To adapt them for the task of MCQ answering, we use the instruction prompt in Figure 2 based on (Plaut et al., 2024) and (Zotos et al., 2024).
@@ 5 Results
Our experiments are aimed at evaluating the usefulness of model uncertainty as a signal for MCQ item difficulty as well as discovering which specific textual and uncertainty features are most relevant for our trained Regressor. We first focus on the performance of our setup using different feature sets (Section 5.1), followed by an in-depth analysis of the importance of individual features (Section 5.2).
Biopsychology BEA
Dummy Regressor 0.1670 0.3110
Best BEA 2024 Competition Result (Team EduTec) - 0.2990
Only Text Features
TF-IDF 0.1479 0.3092
BERT Embeddings 0.1498 0.3066
Only Model Uncertainty Features
1st Token Probabilities 0.1539 0.2960
Choice Order Sensitivity 0.1582 0.3178
1st Token Probability & Choice Order Sensitivity 0.1538 0.2968
Text and Model Uncertainty Features
TF-IDF BERT TF-IDF BERT
First Token Probability 0.1365 0.1385 0.2851 0.2854
Choice Order Sensitivity 0.1309 0.1411 0.2951 0.2961
1st Token Probability & Choice Order Sensitivity 0.1371 0.1388 0.2856 0.2846
Table 3: Root Mean Squared Error (RMSE, the lower the better) on the test set using different sets of features. Lowest achieved RMSE per dataset is shown in boldface. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
@@ 5.1 Performance on Difficulty Estimation
To evaluate the performance of our trained models we use the Root Mean Squared Error (RMSE) metric from Python’s Scikit-learn library Pedregosa et al. (2011), as used in the BEA 2024 Shared task. As previously mentioned, we use a Random Forest Regressor tasked to predict the difficulty of a question item, given as input a vectorised representation of the MCQ as well as the uncertainty of multiple LLMs answering the same MCQ. This creates a modular setup that allows easy manipulation of the input feature set. We present the feature sets along with their performance on the two datasets in Table 3. For brevity, we report the results obtained using Bio-Clinical BERT Embeddings in Appendix D, as they were found to lead to similar, yet consistently slightly worse RMSE scores compared to the general BERT embeddings.
@@
An important first observation is that the RMSE difference between experiments is minimal. This is in-line with the findings from the BEA 2024 shared task, where the lowest achieved RMSE was only $0.0120.012$ lower than the baseline. However, even though the margins are narrow, there are consistent differences between the experimental setups. Most importantly for this research, it is clear that using the uncertainty of the models, combined with text features yields significantly lower RMSE for both datasets, even beating the best score achieved during the BEA competition by a margin of $0.01250.0125$. In fact, providing the Random Forest Regressor only with LLMs’ uncertainties also surpasses the best BEA competition result, albeit by a narrower margin of $0.0030.003$, further underscoring the potential of using model uncertainty for this task. Furthermore, in our exploration we did not find a consistent superiority for one of the two model uncertainty metrics (1st Token Probability or Choice Order Sensitivity) or text vectorisation methods (TF-IDF or BERT embeddings), though there seemed to be a general advantage in using TF-IDF scores over BERT Embeddings. Still, the observed scores for this subset of results are within the observed standard deviation of $0.0020.002$ between repetitions. Lastly, the usefulness of the model uncertainty features is especially clear in the experiments where the Random Regressor did not use any text features: except where only Choice Order Sensitivity is used in the BEA dataset, the performance is consistently better than the respective Mean Regressor baselines.
Figure 3: Biopsychology Dataset. SHAP summary plot showing the contribution of the top ten uni/bi-gram features to the Random Forest’s predictions, highlighting their importance and impact direction. Features are ranked by their average influence, with dots representing individual question items and colour indicating TF-IDF scores. Results averaged over ten repetitions.
5.2 Importance of Different Features
@@
In order to better understand which features drive the predictions of the Random Forest Regressor, we use Shapley additive explanations as provided by the SHAP Python library Lundberg and Lee (2017). To maintain conciseness, we present SHAP summary plots for a selected subset of experiments that we found to be the most insightful. Additional analyses are presented in Appendix E.
Before exploring the analysis regarding model uncertainty, we examine the contribution of the most impactful uni/bi-grams from the text-only experiment using the Biopsychology dataset.
This analysis relies on TF-IDF scores, as BERT embeddings cannot directly be traced back to individual words. Figure 3 shows the ten most impactful features, along with their effect on the Regressors’ prediction for each MCQ item. High TF-IDF scores, highlighted in red, indicate that an MCQ item is predicted to be more difficult (i.e., fewer students answer it correctly). For instance, questions containing the unigram "visual" prompted the Random Forest model to predict greater difficulty. Interestingly, this analysis also demonstrates that questions where a gap (represented by an underscore "_") needs to be filled (e.g., "fill-the-gap" or sentence completion) are predicted to be easier.
(a) Biopsychology Dataset
(b) BEA Dataset
@@
Figure 4: SHAP summary plots for the Biopsychology and BEA datasets showing the contribution of the top ten features to the Random Forest’s predictions. Higher First Token Probability and Order Probability metrics indicate greater model certainty. Results averaged over ten repetitions.
@@
While this analysis gives some insights on important text features, we are mostly interested in the contribution of features related to model uncertainty. Figures 4(a) and 4(b) present the effect of the most impactful features for the feature sets that lead to the best performance (using TF-IDF scores for the text encoding and model uncertainty) for the Biopsychology and BEA datasets, respectively. In both instances, the Random Forest Regressor heavily relies on model uncertainties to predict item difficulty. As hypothesised, the higher the model certainty (in terms of either 1st Token or Choice-Order Probability) the more students are predicted to answer the question correctly. In terms of models, we observe that the confidence of Qwen2.5-14B-chat is especially useful for both datasets (as can be seen by its 1st and 2nd place in Figures 4(a) and 4(b) respectively). This observation also highlights the core challenge of our approach: having a model that is sufficiently capable of answering the MCQs but not so complex that it answers them with complete confidence. In our work, this challenge is partially addressed using an ensemble of models, leaving it up to the Random Forest Regressor to determine their usefulness.
@@ 6 Discussion and Conclusion
We explored how model uncertainty can be used as a proxy for item-difficulty using two MCQ datasets focusing on factual knowledge. We demonstrate, in a simple experimental setup, that while both textual features (e.g., encoding using TF-IDF Scores or BERT embeddings) and model uncertainty features are useful for the task, the trained Random Forest Regressor performed significantly better when model uncertainty features were included.
Our results suggest that aspects of a question item that challenge students similarly impact LLMs. A factor that could explain this alignment is representation: Knowledge that is well represented in an LLM’s training data is likely to be more foundational (e.g., “What is a neuron"), compared to specialised knowledge (e.g., a medical diagnosis). By extension, using model uncertainty for this task requires a model of appropriate size/capabilities.
Our methodological design is intentionally simple, serving as a proof of concept for this approach. This simplicity stems from various design choices. Firstly, our definition of item difficulty is simply the number of students answering an MCQ correctly. This is in contrast to finer-grained metrics such as Item Response Theory Lord and Novick (2008) which however require individual students’ responses that are unavailable in our datasets. Secondly, we use a variety of LLMs without placing great emphasis on their uncertainty behaviour. Specifically, while we ensure that the measured model uncertainty aligns with model correctness (as shown in Appendix A), we do not focus on calibrating the LLMs. Instead, we rely on the Random Forest Regressor to select and weight the uncertainties of the various models. Thirdly, we limited this research to only the necessary features for the purpose of our study (assessing the contribution of model uncertainty), namely text and model uncertainty. In order for yet further improvements to be obtained on the actual task of difficulty estimation, more complex features can be explored and incorporate. For example, building on the intuition that the nature of the distractors plays a role in the question’s difficulty, we experimented with a choice similarity metric, defined (per MCQ) as the average cosine similarity between each distractor and the correct answer choice, similar to the approach by (Susanti et al., 2020). Incorporating this information in our model yields an RMSE of 0.2841 on the BEA dataset, the lowest observed sofar. This improvement is only observed when this feature is combined with model uncertainty. [^8] Experimental details and full results for this additional feature are presented in Table 8 in Appendix F. Shifting into a broader perspective, our findings incentivise further research into better understanding which methods capturing model uncertainty, ranging from analysing model internals to studying model output (e.g., related to prompt sensitivity) are most beneficial for this task.
@@
Indisputably, the central limitation of our approach is the reliance on (un)certain LLMs. As seen in Section 5, model uncertainty is beneficial only when the model can answer the question without being overly confident. As a result, limits the usefulness of our approach, especially given the rapid development of LLMs in terms of their capabilities. We hypothesise that this limitation can at least partially be resolved by using calibrated LLMs, which we leave for future work.
Similarly, our approach is not expected to perform as well on MCQs designed to test knowledge at lower education levels (e.g., primary school geography exams), as even small LLMs are now capable of confidently answering such questions. At the same time, using less proficient LLMs introduces different challenges, particularly regarding linguistic ability: Smaller LLMs are more strongly affected by linguistic perturbations (e.g., question formulation, choice order) and have greater limitations in instruction-following capabilities Biderman et al. (2024); Sclar et al. (2023).
Lastly, due to dataset availability, we evaluated our approach solely on factual knowledge examinations. It remains unclear whether model uncertainty could also be beneficial for assessing the difficulty of examinations of other skill sets, such as language comprehension or mathematical reasoning.
Ethics Statement
In this study, we used a dataset of multiple-choice questions from the "Biopsychology" course at the Behavioural and Social Sciences Faculty of our institution. The data was aggregated across multiple students and anonymised, ensuring that individual student performance cannot be traced.

### v2

@@ 4.2 Model Uncertainty
The current methodological approach is founded on the premise that model uncertainty correlates with student performance and thus, by extension, offers a useful signal when estimating the difficulty of a question item. To explore this hypothesis, we have conducted experiments using two metrics that are shown to correlate well with model correctness (as discussed in Appendix  A): 1st Token Probability and Choice Order Sensitivity. These uncertainty scores are obtained for each LLM separately and concatenated into a single vector, to which any additionally textual, linguistic or choice similarity features are (optionally) also added. This vector is then fed to the regressor.
@@
The first technique to measure model uncertainty is by inspecting the softmax probability of the 1st token to be generated as the answer id to the given MCQ question, (e.g., probability of generating token “B"), in comparison to the probabilities of the alternatives (e.g., probability of generating token “A" or “C"). As the 1st token probabilities can be influenced by the order in which the choices are provided in the problem set [33, 31, 32, 38], we create ten random different orderings for each question and let the model answer each MCQ ten times [^4] For questions with only 3 choices, we instead consider all six different choice orderings.. This way, we calculate the average probability per MCQ choice. We then consider the average probability for the correct answer as the uncertainty metric of the LLM.
Furthermore, as different tokens might be generated to represent the same answer (e.g., “A", “ A", “a ", see details on prompting and answer elicitation in Section 4.4 below) and different models might attribute higher likelihood for specific tokens, the token representing each choice with the highest probability is selected. For example if for a given model the probability of generating token “C" is higher than the probability of token “c", the former is considered for that model. Lastly, the three extracted mean probabilities of all orderings are normalised in the range of $0$ to $11$.
@@
Pezeshkpour and Hruschka [21] observed that choice order sensitivity correlates with error rate. In other words, when LLMs consistently select a choice regardless of its position, that choice is more likely to be correct. Based on this observation, we leverage this correlation to measure uncertainty. Specifically, for all evaluated choice orderings, we measure the probability of the correct choice being selected. Thus, this probability is not based on token probabilities but rather on the eventual choice.
4.3 Additional Features
We also conduct experiments with two additional feature sets: linguistic and choice similarity features. First, based on the work of Ha et al. [8], we extract 17 higher-level linguistic features from each question item. These features vary in complexity, ranging from the number of sentences to the occurrence of additive connectives. Moreover, we also develop a simple baseline that uses these linguistic features along with word2vec embeddings [14], similar to the work by Yaneva et al. [35] [^5] It is worth noting that in contrast to Yaneva et al. [35], we do not use readability metrics as some question items do not meet the $100100$ word requirement for these metrics to be computed.
In the current work we also explore the use of choice similarity defined (per MCQ) as the average cosine similarity between each distractor and the correct answer choice, similar to the approach by [28]. This is operationalised using the Sentence Transformer library [25] with one of two models: all-MiniLM-L6-v2666https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (general embeddings) and S-PubMedBert-MS-MARCO [^7] https://huggingface.co/pritamdeka/S-PubMedBert-MS-MARCO (medical/health text domain embeddings). The two setups are henceforth referred to as “General Similarity" and “Medical Similarity" respectively.
@@ 4.4 Choice of Models and Prompting
In this work, we focus on Decoder-Only models, as they are considered to have incorporated greater amounts of factual knowledge as a byproduct of their language modeling objective [37], compared to Encoder-Only or Encoder-Decoder models. Moreover, as the internal logit probabilities of the 1st token to be generated are needed to measure the uncertainty of each model, we focus on nine open-sourced models of different parameter sizes and families. Additionally, we use the models’ instruction-tuned variant and experiment with both default precision and 4-bit quantised models. This comparison is important because, although quantized models are more efficient (albeit slightly less capable), it is unclear whether they are also more miscalibrated [^8] Details of the models used are in Table 6 in Appendix B.. To adapt them for the task of MCQ answering, we use the instruction prompt in Figure 3 based on the experimental setup of Plaut et al. [22] as well as Zotos et al. [40].
[…]
Instruction phrasing used for all models and experiments: Below is a multiple-choice question. Choose the letter which best answers the question. Keep your response as brief as possible; just state the letter corresponding to your answer with no explanation.
@@
Table 3: Root Mean Squared Error (RMSE, the lower the better) on the test set using different sets of features. Lowest achieved RMSE per dataset is shown in boldface. Best overall results are highlighted in light green. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
Method Biopsychology USMLE CMCQRD CMCQRD_IRT
Baselines
Dummy Regressor 0.1667 0.3110 0.1833 9.7568
Best Literature Result [24] - 0.291 - 8.5
Only Text
TF-IDF 0.1479 0.3092 0.1843 9.2531
BERT Embeddings 0.1498 0.3066 0.1843 8.9347
Only Uncertainty
1st Token Probabilities 0.1473 0.3041 0.1770 9.4161
Choice Order Sensitivity 0.1543 0.3155 0.1788 9.9787
Both Uncertainty Features 0.1460 0.3034 0.1754 9.3705
Text and Uncertainty
TF-IDF BERT TF-IDF BERT TF-IDF BERT TF-IDF BERT
First Token Probability 0.1361 0.1362 0.3037 0.2868 0.1668 0.1669 8.5661 8.2763
Choice Order Sensitivity 0.1378 0.1409 0.310 0.2906 0.1658 0.1676 8.6663 8.4204
Both Uncertainty Features 0.1359 0.1365 0.3044 0.2864 0.1654 0.1669 8.5933 8.3009
@@ 5 Results
Our experiments are aimed at evaluating the usefulness of model uncertainty as a signal for MCQ item difficulty as well as discovering which specific textual and uncertainty features are most relevant for our trained Regressor. We first focus on the performance of our setup using text and model uncertainty feature sets (Section 5.1), followed by an exploration of the effect of additional features (Section 5.2). We conclude with a post-hoc analysis of our system using SHAP explanations (Section 5.3). While the following sections present only the relevant results, Appendix C provides an overview of the results for all experiments. All experiments were conducted using two Nvidia A100 GPUs.
@@ 5.1 Performance on Difficulty Estimation
To evaluate the performance of our trained models we use the Root Mean Squared Error (RMSE) metric from Python’s Scikit-learn library [18], as used in the BEA 2024 Shared task. As previously mentioned, we use a Random Forest Regressor tasked to predict the difficulty of a question item, given as input a vectorised representation of the MCQ as well as the uncertainty of multiple LLMs answering the same MCQ. This creates a modular setup that allows easy manipulation of the input feature set. We present the feature sets along with their performance on the three datasets in Table 3. For brevity, we report the results obtained using Bio-Clinical BERT Embeddings in Appendix C, as they were found to lead to similar, yet consistently slightly worse RMSE scores compared to the general BERT embeddings. Similarly, Table 3 presents the performance of the default precision variants of the models, while the results for the quantised models are reported in Appendix C. Lastly, for CMCQRD, we report the performance on predicting both p-values and the IRT-based metric.
@@
An important first observation is that the RMSE difference between experiments is minimal. This is in-line with the findings from the BEA 2024 shared task, where the lowest achieved RMSE was only $0.0120.012$ lower than the baseline, and the achieved RMSE scores of the top 10 approaches were within 0.009. Even so, there are consistent differences between the experimental setups. Most importantly for this research, incorporating model uncertainty alongside text features significantly reduces RMSE across all datasets, outperforming the best scores in previous literature. Even lower RMSE is achieved for the USMLE and CMCQRD datasets when additional features are included (see Section 5.2). Furthermore, our exploration revealed that 1st Token Probability consistently serves as a more useful signal for the task than Choice Order Sensitivity. However, combining both yields the best results. Regarding the two text vectorization methods, we find no significant differences between them, except for the USMLE dataset, where BERT embeddings outperform TF-IDF scores.
5.2 Effect of Additional Features
Naturally, additional features of an MCQ might capture aspects that make the question easier or more challenging. Table 4 explores the effect of higher-level linguistic features and the similarity between choices (as described in Section 4.3). Expectedly, using choice similarity as the only feature yields poor results on the task. Similarly, the baseline that focuses on linguistic features only marginally beats the Mean Regressor baseline for two of the four experiments. However, combining either, or both, of these features with text and model uncertainty features further improves our best results for all but the Biopsychology experiments (where the best result is still achieved without the additional features). Moreover, we observe that there is no clear advantage between the General or Medical Similarity, with the latter also seemingly being useful in the non-medical domain (CMCQRD dataset). These experiments highlight that while model uncertainty and basic text encodings (such as BERT or TF-IDF scores) capture certain aspects of difficulty, other factors should still be utilized for this task.
Table 4: Root Mean Squared Error (RMSE, the lower the better) on the test set using additional features. Lowest achieved RMSE per dataset is shown in boldface. Best overall results, also in comparison to those presented in Table 3, are additionally highlighted in light green. All results are averaged over ten repetitions, with the standard deviation not exceeding 0.002.
Method Biopsychology USMLE CMCQRD CMCQRD_IRT
Baselines
Dummy Regressor 0.1667 0.3110 0.1833 9.7568
Best Literature Result [24] - 0.291 - 8.5
Linguistic Features Baseline [8] 0.1544 0.3147 0.1852 9.3335
Only Choice Similarity
General (all-MiniLM-L6-v2) 0.1895 0.3567 0.2226 12.2829
Medical (S-PubMedBert-MS-MARCO) 0.1883 0.3432 0.2146 11.4768
TF-IDF BERT TF-IDF BERT TF-IDF BERT TF-IDF BERT
Text, Both Uncertainties & Choice Similarity
General Similarity 0.1372 0.1367 0.2835 0.2862 0.1651 0.1669 8.5934 8.2956
Medical Similarity 0.1376 0.1361 0.2836 0.2853 0.1651 0.1664 8.5999 8.2325
Both Similarities 0.1378 0.1365 0.2847 0.2862 0.1653 0.1676 8.5694 8.2817
Text, Both Uncertainties, Both Sim & Linguistic Features
0.1393 0.1362 0.2817 0.2857 0.1652 0.1673 8.5490 8.6575
5.3 Feature Importance
@@
In order to better understand which features drive the predictions of the Random Forest Regressor, we use Shapley additive explanations as provided by the SHAP Python library [13]. To maintain conciseness, we present SHAP summary plots for a selected subset of experiments that we found to be the most insightful.
[…]
SHAP summary plot showing the contribution of the top ten uni/bi-gram features to the Random Forest’s predictions Refer to caption
Figure 4: Biopsychology Dataset. Shapley summary plot showing the contribution of the top ten uni/bi-gram features to the Random Forest’s predictions, highlighting their importance and impact direction. Features are ranked by their average influence, with dots representing individual question items and colour indicating TF-IDF scores. Results averaged over ten repetitions.
[…]
Model Uncertainty has a high effect on prediction. Some words are also influential, but the most influence is from a variety of models
(a) Biopsychology
(b) USMLE
(c) CMCQRD (p-value)
@@
Figure 5: Shapley summary plots for the three datasets showing the contribution of the top ten features to the Random Forest’s predictions. Higher First Token Probability and Order Probability metrics indicate greater model certainty. Results averaged over ten repetitions.
Before exploring the analysis regarding model uncertainty, we examine the contribution of the most impactful uni/bi-grams from the text-only experiment using the Biopsychology dataset. This is useful because it allows us to gain an overview of the influence of lexical features before introducing model uncertainty, while also highlighting any features that unexpectedly influence question difficulty (e.g., interrogative words).
This analysis relies on TF-IDF scores, as BERT embeddings cannot directly be traced back to individual words. Figure 4 shows the ten most impactful features, along with their effect on the Regressors’ prediction for each MCQ item. High TF-iDF scores are highlighted in red and broadly represent the presence of the word in an item. Furthermore, the impact on the Regressor’s prediction is either positive or negative, meaning that a feature can either lower or increase the predicted difficulty score. As can be seen, the presence of certain terms (e.g., “visual", “hippocampus") lead the Regressor to predict higher difficulty. Interestingly, this analysis also demonstrates that questions where a gap (represented by an underscore “_") needs to be filled (e.g., “fill-the-gap" or sentence completion) are predicted to be easier.
While this analysis shows that the presence of certain words can steer the Regressor towards predicting a higher or lower difficulty, we are mostly interested in the contribution of features related to model uncertainty. Figures 5(a),  5(b) and 5(c) present the effect of the most impactful features for the feature sets that lead to the best performance (using TF-IDF scores for the text encoding and model uncertainty) for the Biopsychology, USMLE and CMCQRD datasets, respectively. It is worth noting that for the CMCQRD dataset, we focus on the prediction of the number of students selecting the correct rate (p-value), instead of the IRT-based metric.
@@
In all instances, the Random Forest Regressor heavily relies on model uncertainties to predict item difficulty. As hypothesised, the higher the model certainty (in terms of either 1st Token or Choice-Order Probability) the more students are predicted to answer the question correctly. In each configuration, the uncertainty of different models has the greatest influence. Notably, the uncertainty of Qwen models consistently serves as a strong indicator of difficulty. Furthermore, this analysis hints towards model size being important, especially when comparing the Biopsychology and the USMLE results: For the latter, the confidence of larger models is more influential in the Regressor’s prediction of question difficulty. This observation also highlights the core challenge of our approach: having a model that is sufficiently capable of answering the MCQs but not so complex that it answers them with complete confidence. In our work, this challenge is partially addressed using an ensemble of models, leaving it up to the Random Forest Regressor to determine their usefulness.
@@ 6 Discussion and Conclusion
We explored how model uncertainty can be leveraged for the task of question item difficulty estimation using three MCQ datasets focusing on factual knowledge and language comprehension. We demonstrate, in experimental setups of varying complexity, that while both textual features (e.g., encoding using TF-IDF Scores or BERT embeddings) and model uncertainty features are useful for the task, the trained Random Forest Regressor performed significantly better when model uncertainty features were included.
Our results suggest that aspects of a question item that challenge students similarly impact LLMs. A factor that could explain this alignment is representation: Knowledge that is well represented in an LLM’s training data is likely to be more foundational (e.g., “What is a neuron"), compared to specialised knowledge (e.g., a medical diagnosis). By extension, using model uncertainty for this task requires a model of appropriate size/capabilities. Additionally, our results suggest that an LLM’s uncertainty does not fully account for certain aspects of item difficulty, such as linguistic complexity [^9] Notably, a question item can be complex yet easy, or vice versa [19]..
Our methodological design is intentionally simple, serving as a proof of concept for this approach. This simplicity stems from various design choices. Firstly, we use a variety of LLMs without placing great emphasis on their uncertainty behaviour. Specifically, while we ensure that the measured model uncertainty aligns with model correctness (as shown in Appendix A), we do not focus on calibrating the LLMs. Instead, we rely on the Random Forest Regressor to select and weight the uncertainties of the various models. Secondly, we conducted a series of experiments using features of varying complexity to demonstrate that, while incorporating additional features (e.g., average cosine similarity between each distractor and the correct answer choice) can improve task performance, the improvements are often marginal compared to relying solely on a simple text encoding combined with model uncertainty.
Shifting into a broader perspective, our findings suggest that there are similarities between the way LLMs and students process educational material. While caution is necessary, we see potential for future research to leverage LLMs for student and cohort modeling.
@@
Indisputably, the central limitation of our approach is the reliance on (un)certain LLMs. As seen in Section 5, model uncertainty is beneficial only when the model can answer the question without being overly confident. Naturally, this limits the usefulness of our approach, especially given the rapid development of LLMs in terms of their capabilities. We hypothesise that this limitation can at least partially be resolved by using calibrated LLMs, which we leave for future work.
Similarly, our approach is not expected to perform as well on MCQs designed to test knowledge at lower education levels (e.g., primary school geography exams), as even small LLMs are now capable of confidently answering such questions. At the same time, using less proficient LLMs introduces different challenges, particularly regarding linguistic ability: Smaller LLMs are more strongly affected by linguistic perturbations (e.g., question formulation, choice order) and have greater limitations in instruction-following capabilities [4, 26].
Lastly, due to dataset availability, we evaluated our approach solely on three examinations. It remains unclear whether model uncertainty could also help assess the difficulty of exams in other skill sets, such as mathematical reasoning.